OpenStudy (anonymous):

what is the simplified form of 12y^4/8x^3 times 6x^8/20 y^6

OpenStudy (campbell_st):

it can be written as \[\frac{12y^4 \times 6x^8}{8x^3 \times 20y^6} = \frac{8x^3y^4(9x^5)}{8x^3y^4(20y^2)}\] cancel the common factor for the answer